deidrenicoleg:

Rewrite the expression with rational exponents as a radical expression by extending the properties of integer exponents. three to the one third power all over three to the one sixth power the square root of three to the one sixth power the ninth root of three squared the square root of three to the sixth power the sixth root of three

deidrenicoleg:

I think it is the square root of three to the sixth power am i correct?

TheSmartOne:

In math, there is no thinking Your work should lead you to one answer Do you know the exponent rules?

TheSmartOne:

\(\frac{a^b}{a^c}=a^{b-c}\) \( a^{\frac{b}{c}} = \sqrt[c]{a^b}\)